Please help me with this problem, using the Pythagorean theorem:
A rectangular park is 6 miles long and 3 miles wide. How long is a pedestrian route that runs diagonally across the park?
Please let me know the steps that lead you to the solution. Thanks so much!

Answer by nerdybill(7384)

A rectangular park is 6 miles long and 3 miles wide. How long is a pedestrian route that runs diagonally across the park?
.
the "diagonal" is the hypotenuse so from Pythagorean we have:
d^2 = 6^2 + 3^2
d^2 = 36 + 9
d^2 = 45
d = sqrt(45)
d = sqrt(3*3*5)
d = 3*sqrt(5) miles (exact answer)
.
d = 6.71 miles (approx answer)